Find the critical value based on the given information. Н. 0 < 0.629 n = 19, a = 0.025. 8.907 31.526 O 7.015 8.231
Answer #1

Correct answer: Option (D) 8.231

since H1: Sigma < 0.629 that means, it is left tail test, alpha = 0.025

It is test for single variance and follows chi-square test for single variance, the df is n - 1 = 19-1 = 18

From chi-square tables, The critical value of chi-square is 8.231 at 0.025 in left tail tes and 18 df

Graph:

Distribution Plot Chi-Square, df=18 0.07 0.06 0.05 0.04 Density 0.03 0.02 0.01 0.025 0.00 0 8.231 х